OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? A bird house is in the shape of a square pyramid. If the volume is 180 cubic inches, how tall is the bird house if the area of the base is 9 square inches? a. 60 inches b. 21 inches c. 22 inches d. 23 inches

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

ok so volume of a square pyramid is \(a^2\dfrac{h}{3}\) where a is the base edge. they tell us that 9 is the area of the base not just the edge which means that it has already been squared

OpenStudy (sleepyjess):

following that part of it?

OpenStudy (anonymous):

@sleepyjess yes

OpenStudy (sleepyjess):

sorry i had to restart my computer because the internet wasnt working

OpenStudy (sleepyjess):

ok so now we have \(180=9\dfrac{h}{3}\) since you are multiplying the right side by 9 now you are going to divide both sides by 9 to get rid of 9 which leaves you with \(20=\dfrac{h}{3}\) since you are dividing by 3 what do you think you will do to get rid of 3?

OpenStudy (sleepyjess):

you still here? @ninaesb

OpenStudy (anonymous):

divide 20 by 3 ? @sleepyjess

OpenStudy (sleepyjess):

no you are dividing h by 3 so since you are originally dividing you want to do the opposite

OpenStudy (anonymous):

so divide 3 by h so then it would be 20=3h ?

OpenStudy (sleepyjess):

so now you multiply both sides by 3 leaving you with ?=h

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

no if you multiply 3*h/3 the 3's cancel out so you have 20*3=h

OpenStudy (anonymous):

@sleepyjess so its 60 ?

OpenStudy (sleepyjess):

yes

OpenStudy (anonymous):

thankyou @sleepyjess

OpenStudy (sleepyjess):

if you want you can test it \(9\dfrac{60}{3}\) 9*20 180